I've opened the AP table and I'm working on it.
I re-aligned the Faraday on the AP table. I also aligned the beam to the periscope on the PSL and all the other optics along the beam path. Now I have a nice NPRO beam at the PLL which overlaps with the PSL beam. The alignment has to be further improved becasue I see no beat yet.
I wonder if the all the tinkering on the PSL laser done recently to revive the power have changed the PSL NPRO temperature and so its frequency. That could also explain why the beat doesn't show up at the same temperature of the NPRO as I used to operate it. Although I scanned the NPRO temperature +/- 2 deg and didn't see the beat. So maybe the misalignment is the casue.
Not feeling very well right now. I need to go home for a while.
AP table closed at the moment.
NPRO shutter closed
We definitely changed the PSL NPRO temp while fiddling around, trying to increase the laser power. I think it's noted in the elog both times that it's happened in the last few months (once when Rana, Koji and I worked on it, and then again when it was just Koji), but we opened up the side of the MOPA box so that we could get at (and change) the potentiometer which adjusts the NPRO temp. So you may have to search around for a while.